# VGGT-SLAM 2.0: Real time Dense Feed-forward Scene Reconstruction

### v1

@@ IV-C Loop closures and global optimization
Similar to VGGT-SLAM, we perform image retrieval using SALAD [10] to check for possible loop closures between frames in the current submap and frames in prior submaps (ignoring the most recent prior submap since it will always have overlap with the current submap). We differ from VGGT-SLAM in two key ways. Firstly, instead of adding the retrieved frames into the current submap, we create a small submap of two frames (the loop closure submap shown in Figure 3). This enables greater flexibility as VGGT can be run on the current batch of frames before performing image retrieval. This two frame submap consists of the retrieved frame from a prior submap and the corresponding query frame in the current submap. This two frame loop closure submap is passed to VGGT and Inner submap edges are then connected between the frames of the loop closure submap and their respective overlapping frame, as shown in Figure 3.
@@
Given the loop closure constraints, we run global optimization which follows VGGT-SLAM and is optimized on the $\mathrm{SL}(4)$ manifold using GTSAM [40]. The $\mathrm{SL}(4)$ optimization solver has since been added to the official GTSAM release, making installation simpler with just a pip install.
@@
We can recover projection matricies $\mathbf{P_{i}}$ which transform points from the global frame, $w$, to the camera frame corresponding to image $I_{i}$ as
$\mathbf{P}_{i}=\mathbf{K}_{i}(\mathbf{H}^{w}_{i})^{-1}$ (6)
Note that differently from VGGT-SLAM, since we have defined points to be relative to their respective frames and not the first frame of a submap, the right-hand side of Eq. 6 only has a homography multiplied with a camera calibration (augmented to be $4\times 4$) from VGGT instead of additionally being multiplied by a pose from VGGT. Global poses are found by decomposing $\mathbf{P_{i}}$ into calibration and pose, and global 3D points are recovered using Eq. 1 with $\mathbf{H}^{w}_{i}$.
@@ V-C Loop closure verification evaluation
We demonstrate that our image retrieval verification method which was presented in Section IV-B can help achieve more loop closures while preventing false positives. In Table II we compare the number of loop closures found by running VGGT-SLAM 2.0 on the three scenes of the Clio [39] datasets using a SALAD threshold of 0.80 (which was the value used in  [5]) without using retrieval verification to relaxing the value to 0.95 with using retrieval verification. The office scene without verification led to false positive loop closures causing the reconstruction to diverge even with a SALAD threshold of 0.80 due to challenging similar looking areas such as the example from Fig. 5(c). Meanwhile, using verification led to more loop closures and no false positives, even on the office scene.
@@ V-E Additional qualitative results
While we design VGGT-SLAM 2.0 primarily for indoor datasets or small scale outdoor scenes, we also include an example from the Kitti [51] dataset in Fig. 9 to provide an example of running on long outdoor sequences to demonstrate improved robustness to long sequences. VGGT-SLAM diverges on the Kitti dataset due to long distances between look closures causing rapid drift, and vulnerability to planar degeneracies as the majority of VGGT reconstructed points are on a flat road. Both of these scenes (which contain 34 submaps for the barn and 44 for Kitti) are larger than the largest scene shown in [5] which contained only 22 submaps of size 16 frames. Additional visual results are provided in the supplementary video  [^2] https://www.youtube.com/watch?v=GBdOvd6p4OU.
@@
V-F Real time experiments and timing results
The runtime of VGGT-SLAM 2.0 is about 120 ms per frame (8.4 FPS) for a submap of size 16 frames using a 3090 GPU if no open-set CLIP vectors are computed and 158 ms if open-set object detection is enabled. A breakdown of the major timing components of VGGT-SLAM 2.0 is shown in Table IV. Over half of the time per frame is used to run VGGT (78 ms) which means further reduction in computation can benefit from future faster variants of VGGT. Timing is computed by summing the total time during a submap’s construction and dividing by the number of unique frames in the submap (which does not include the overlapping frames). Since timing is very hardware specific, we also run VGGT-SLAM and MASt3R-SLAM on the same machine to provide an approximate reference point for comparison. We observe that MASt3R-SLAM runs at a comparable 7.2 FPS and VGGT-SLAM runs at 6.9 FPS.
Stage time per keyframe (ms)
VGGT Inference 78 ms
Keyframe Detection 11 ms
Loop Closure Detection 8 ms
Backend Optimization 0.5 ms
Semantics (optional) 23 ms
TABLE IV: Average runtime per frame in milliseconds of the major components of VGGT-SLAM 2.0. Timing results reported using a 3090 GPU for submap size of 16 frames.
@@ VI Limitations
While our proposed approach substantially improves VGGT-SLAM and other state-of-the-art methods, it is not without limitations. Through experiments on a variety of custom scenes, we did observe some cases of unsatisfactory results and have includes examples in the supplementary video to provide both a holistic evaluation and present possible directions of future improvements. One such experiment is the reconstruction of an empty four-bedroom house which contained keyframes which only saw a plain white wall causing VGGT reconstruction to fail. In this case, VGGT-SLAM 2.0 reconstruction will also diverge; possible future work could incorporate a lost-tracking module to recover from this. Additionally, our backend factor-graph optimization is only over poses and not points. While this greatly simplifies the SLAM pipeline (such as by not requiring tracking correspondences) and is effective in many of our experiments, in some cases it can cause artifacts from misaligned 3D points and in some challenging scenes lead to potentially preventable failure modes. Finally, the first frame of the submap is also unequally influential since not only does VGGT reconstruct all frames w.r.t. to first frame, but we also use the first frame to compute the scale factor of a submap. Possible future work could optimally select the first frame to both improve scale estimation and reduce VGGT failure modes.

### v2

@@ IV-C Loop closures and global optimization
Similar to VGGT-SLAM, we perform image retrieval using SALAD [10] to check for possible loop closures between frames in the current submap and frames in prior submaps (ignoring the most recent prior submap since it will always have overlap with the current submap). We differ from VGGT-SLAM in two key ways. Firstly, instead of adding the retrieved frames into the current submap, we create a small submap of two frames (the loop closure submap shown in Figure 3). This enables greater flexibility as VGGT can be run on the current batch of frames before performing image retrieval. This two frame submap consists of the retrieved frame from a prior submap and the corresponding query frame in the current submap. This two-frame loop closure submap is passed to VGGT and an Inter submap edge is then added between the frame of the loop closure submap and its respective overlapping frame, as shown in Figure 3.
@@
Given the loop closure constraints, we run global optimization which follows VGGT-SLAM and is optimized on the $\mathrm{SL}(4)$ manifold using GTSAM [41] without robust losses. The $\mathrm{SL}(4)$ optimization solver has since been added to the official GTSAM release, making installation simpler with just a pip install.
@@
We can recover projection matrices $\mathbf{P_{i}}$ which transform points from the global frame, $w$, to the camera frame corresponding to image $I_{i}$ as
$\mathbf{P}_{i}=\mathbf{K}_{i}^{3\times 4}(\mathbf{H}^{w}_{i})^{-1}$ (7)
Note that differently from VGGT-SLAM, since we have defined points to be relative to their respective frames and not the first frame of a submap, the right-hand side of Eq. 7 only has a homography multiplied with a camera calibration (padded with a zeros vector to be $3\times 4$) from VGGT instead of additionally being multiplied by a pose from VGGT. Global poses are found by decomposing $\mathbf{P_{i}}$ into calibration and pose, and global 3D points are recovered using Eq. 1 with $\mathbf{H}^{w}_{i}$.
@@ V-C Loop closure verification evaluation
We demonstrate that our image retrieval verification method which was presented in Section IV-B can help achieve more loop closures while preventing false positives. In Table II we compare the number of loop closures found by running VGGT-SLAM 2.0 on the three scenes of the Clio [40] datasets using a SALAD threshold of 0.80 (which was the value used in [5]) without using retrieval verification against the case where we relax the value to 0.95 and use the proposed retrieval verification. The office scene without verification led to false positive loop closures causing the reconstruction to diverge even with a SALAD threshold of 0.80 due to challenging similar looking areas such as the example from Fig. 5(c). Meanwhile, using verification led to more loop closures and no false positives, even on the office scene.
@@ V-E Additional qualitative results
While we design VGGT-SLAM 2.0 primarily for indoor datasets or small scale outdoor scenes, we also include an example from the Kitti [52] dataset in Fig. 9 to provide an example of running on long outdoor sequences to demonstrate improved robustness to long sequences. VGGT-SLAM diverges on the Kitti dataset due to long distances between loop closures causing rapid drift, and vulnerability to planar degeneracies as the majority of VGGT reconstructed points are on a flat road. Both of these scenes (which contain 34 submaps for the barn and 44 for Kitti) are larger than the largest scene shown in [5] which contained only 22 submaps of size 16 frames. Additional visual results are provided in the supplementary video. [^2] https://www.youtube.com/watch?v=GBdOvd6p4OU
@@
V-F Real-time experiments and timing results
The runtime of VGGT-SLAM 2.0 is about 8.4 FPS for a submap of size 16 frames using a 3090 GPU if no open-set CLIP vectors are computed, and 6.3 FPS if open-set object detection is enabled. FPS is computed by diving the total time to handle a submap by the number of unique frames in the submap (which does not include the overlapping frames). A breakdown of the major timing components of VGGT-SLAM 2.0 is shown in Table IV. Over half of the time per submap is used to run VGGT (1248 ms) which means further reduction in computation can benefit from future faster variants of VGGT. Since timing is very hardware specific, we also run VGGT-SLAM and MASt3R-SLAM on the same machine to provide an approximate reference point for comparison. We observe that MASt3R-SLAM runs at a comparable 7.2 FPS and VGGT-SLAM runs at 6.9 FPS.
Stage total time per submap (ms)
VGGT Inference 1248 ms
Keyframe Detection 176 ms
Loop Closure Detection 128 ms
Backend Optimization 8 ms
Semantics (optional) 368 ms
TABLE IV: Average runtime per submap in milliseconds of the major components of VGGT-SLAM 2.0. Timing results reported using a 3090 GPU for submap size of 16 frames.
@@ VI Limitations
While our proposed approach substantially improves VGGT-SLAM and other state-of-the-art methods, it is not without limitations. Through experiments on a variety of custom scenes, we did observe some cases of unsatisfactory results and have includes examples in the supplementary video to provide both a holistic evaluation and present possible directions of future improvements. One such experiment is the reconstruction of an empty four-bedroom house which contained keyframes which only saw a plain white wall causing VGGT reconstruction to fail. In this case, VGGT-SLAM 2.0 reconstruction will also diverge; possible future work could incorporate a lost-tracking module to recover from this. Additionally, our backend factor-graph optimization is only over poses and not points. While this greatly simplifies the SLAM pipeline (such as by not requiring tracking correspondences) and is effective in many of our experiments, in some cases it can cause artifacts from misaligned 3D points and in some challenging scenes leads to potentially preventable failure modes. Finally, the first frame of the submap is also unequally influential since not only does VGGT reconstruct all frames w.r.t. to first frame, but we also use the first frame to compute the scale factor of a submap. Possible future work could optimally select the first frame to both improve scale estimation and reduce VGGT failure modes.